OpenStudy (anonymous):

help with 13+7=32-x and show your work

OpenStudy (balala):

Alrighty, so the first thing you can do is simplify a little bit :) Just add the 13 and 7 together... 20 = 32 - x Now we isolate 'x' 20 = 32 - x -32 - 32 Now your equation should look like this: -12 = -x Now are last step is to divide both sides of the equation by -1 :) -12 = -x /-1 /-1 12 = x and ta da! x is equal to 12! :)